Large Consignments of computer components are inspected for
defectives by means of a sampling system. Ten components are
examined and the lot is to be rejected if two or more are found
to be defective. If a consignments contains exactly 10%
defectives.
Find the probability of the consignment by using the technique
of Binomial probability distribution that the consignment is:
i) Accepted
ii) Rejected
1
Expert's answer
2021-07-19T05:46:05-0400

Let "X=" the number of defective components: "X\sim Bin(n, p)."

Given "n=10, p=0.1, q=1-p=1-0.1=0.9."

i)


"P(X<2)=P(X=0)+P(X=1)"

"=\dbinom{10}{0}(0.1)^0(0.9)^{10-0}+\dbinom{10}{1}(0.1)^1(0.9)^{10-1}"

"=0.7360989291"

The probability that the consignment is accepted is 0.7360989291.



ii)


"P(X\geq2)=1-P(X<2)"

"=1-\dbinom{10}{0}(0.1)^0(0.9)^{10-0}-\dbinom{10}{1}(0.1)^1(0.9)^{10-1}"

"=0.2639010709"

The probability that the consignment is rejected is 0.2639010709.